OpenStudy (anonymous):

Find the solutions for x^2+1=0.

OpenStudy (anonymous):

\[x^2=-1\] is a start

OpenStudy (anonymous):

\[x^2 +1=0 \] \[x^2=-1\] \[x=i\]

OpenStudy (anonymous):

Is the solution just i or -i as well?

OpenStudy (anonymous):

\[i=\sqrt{-1}\] so -i would actually be \[-\sqrt{-1}\]

OpenStudy (loser66):

@idealist hey girl, this question from which part? complex matrix? be careful with it, it links to Euler's formula for the solution of ODE

OpenStudy (anonymous):

Yes, I will.

OpenStudy (loser66):

ok